Volume of the Solid (Shell Method)


Given the bounded region is revolved about y axis, find the volume of the solid generated.


The region below the curve y= Inx, above the x-axis, and to the left of the line x=4.


1
Expert's answer
2022-05-30T03:56:45-0400
"V=2\pi\displaystyle\int_{0}^4x\ln xdx"

"\int x\ln x dx"

"u=\ln x, du=\dfrac{dx}{x}"

"dv=xdx, v=\dfrac{x^2}{2}"

"\int x\ln x dx=\dfrac{x^2\ln x}{2}-\dfrac{1}{2}\int xdx"

"=\dfrac{x^2\ln x}{2}-\dfrac{x^2}{4}+C"

"V=2\pi\displaystyle\int_{0}^4x\ln xdx"

"=2\pi\lim\limits_{t\to 0^+}\displaystyle\int_{t}^4x\ln xdx"

"=2\pi\lim\limits_{t\to 0^+}[\dfrac{x^2\ln x}{2}-\dfrac{x^2}{4}]\begin{matrix}
 4 \\
 t
\end{matrix}"

"=2\pi(\dfrac{4^2\ln 4}{42}-\dfrac{4^2}{4}-(0-0))"

"=8\pi(4\ln 2-1)\ ({units^3})"